A 1000 kg car is on a 25 cm piston. If your mass is 60 kg, what size of piston would you need to lift the 1000 kg car by standing on the piston?
1
Expert's answer
2020-02-21T10:01:02-0500
"p_1=p_2"

"\frac{Mg}{0.25\pi d^2}=\frac{mg}{0.25\pi d'^2}\to \frac{M}{d^2}=\frac{m}{ d'^2}"

"\frac{1000}{25^2}=\frac{60}{ d'^2}"

"d'=6.1\ cm"